8. A university bookstore recently sold a wirebound graph-paper notebook for $2.55, and a college-ruled notebook for $3.22. At the start of spring semester, a combination of 50 of these notebooks were sold for a total of $157.65. How many of each type were sold?

thanks for any help!!!

Answer by nerdybill(7384)

8. A university bookstore recently sold a wirebound graph-paper notebook for $2.55, and a college-ruled notebook for $3.22. At the start of spring semester, a combination of 50 of these notebooks were sold for a total of $157.65. How many of each type were sold?
.
Let x = number of graph-paper notebooks sold
then
50-x = number of college-ruled notebooks sold
.
2.55x + 3.22(50-x) = 157.65
2.55x + 161-3.22x = 157.65
161-0.67x = 157.65
-0.67x = -3.35
x = 5 (graph-paper notebooks)
.
college-ruled notebooks:
50-x = 50-5 = 45